Using a five-sided spinner (red, yellow, green, teal, blue) you construct or a digital five-sided spinner, complete the following steps; write the probabilities as a fraction and a percentage.
1. What is the theoretical probability of spinning yellow?
2. What is the theoretical probability of spinning red or yellow?
3. Spin the spinner 20 times, create a table to display the results. What is the experimental probability of spinning yellow?
4. Perform another 20 spins, and record them. What is the new experimental probability of spinning yellow out of all 40 spins?
5. What is the experimental probability of spinning red or yellow out of all 40 spins?

Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52787)

.

Hey, I read your n.n. 3 and 4, where you request the tutors to spin the spinner 20 times
for you and then to spin it 20 more times, and I ask myself (and ask you, too, very politely) :
are you OK posting such requests to tutors?

These requests seem very strange to me . . .



Answer by math_tutor2020(3817)


Problem 1

Answer: 1/5 = 20%

Reason: There is 1 section we want (yellow) out of 5 sections total.
This assumes each section is the same area, which leads to each section having equal chance of being landed on.

=======================================================

Problem 2

Answer: 2/5 = 40%

Reason: This time there are two sections we want (red and yellow) out of 5 total.

=======================================================

Problem 3 through 5

I'll let the student do this part.
Create a physical spinner that you can hold, or go to any website that will provide a digital spinner.
A random number generator will work as well.

The results of each trial will be different between students. This would mean the results of problems 3 and 4 will be different.

Experimental probability, aka empirical probability, is when we divide the number of observed cases over the total number of trials.

For example, if you see red come up 5 times out of 40 spins total, then the empirical probability of red would be 5/40 = 1/8 = 0.125 = 12.5%
This is a hypothetical example and not the answer to questions 3 through 5.

As the name implies, theoretical probability is based on the theory of each space having equal area (and hence each space has equal chance of being landed on).
In contrast, experimental probability requires the student to do trials/experiments to generate the probability distribution table.

Further Reading
https://www.statology.org/theoretical-probability/